If n(A) = 40, N(B) = 117, and n(AUB) = 137, what is n(A ∩B)?
Answer by MathLover1(20849)

n(A ∩B) = n(A) + n(B) - n(AUB)
n(A ∩B) = 40 + 117 - 137 = 20